### v1

@@ Acknowledgements.
However, it is important to note that the effectiveness of an AES system hinges on the underlying algorithms and models. There has been significant research using different model architectures in the AES task, and although transformer-based models have shown remarkable success in various natural language processing tasks [26], their performance in AES tasks has revealed certain limitations. This was demonstrated by Tashu et al. [39], in contrast to the limitations discussed by Wang et al. [44]. By using a Convolution Neural Network (CNN) and Recurrent Neural Network (RNN) based architecture, Tashu’s DeLAES model significantly outperformed the transformer-based state-of-the-art model of Yang et al. [46]. It is important to note that none of these transformer-based models implemented proper contextual support to leverage the data fully. This inadequacy sparks the need to explore context augmentation techniques that could potentially enhance the capabilities of transformer-based AES models.
Contextual cues have been found to impact human understanding of language significantly [11], and there is a growing recognition of their potential in improving AI-driven language tasks[14]. Our research aims to harness this potential by investigating the integration of context in the context of AES. At the core of this research lies a fundamental question:
@@
We assert that an essay can be described as a textual composition crafted to persuade the reader of the author’s perspective. Consequently, essays primarily strive to argue for a specific standpoint concerning the topic at hand. We substantiate this based on Mercier and Sperber’s research [25], which underscores the intrinsic argumentative nature of human reasoning, further reinforcing the inherently persuasive quality of essays.
To restrict our scope of research, we limited our search to the following granularities: Relative Context, Prompt Context, Structural Context and Feature-based Context. The essays are graded on a curve, implying that an individual essay would have a relative rank amongst all the other essays for a specific prompt. Margin Ranking, as discussed in Section 4.4, explores this in the form of relative context. As discussed in Section 3.1, the dataset has essays collected from eight different sets where every essay is written in response to their respective prompt. Understanding the prompt or the topic should be crucial to assess the essay adequately. This should give the model an insight into the essay’s key elements, ideally better identifying unique patterns for good and bad aspects of a given prompt. Thus, a prompt can successfully act as a proper context. Essays may follow different genres and writing styles, such as argumentative, narrative, or expository. However, we assume that an essay’s primary objective is to convince the reader of some truth. To this end, if we provide the latent structural context of the essay to the model, it may understand the interdependencies of the sentences better. To address this, we want to introduce the structural contexts at the scale of Elementary Discourse Units (EDUs) [24] and Argument Components (ACs) [20]. As discussed in Section 2.2, the length is correlated with the scores in the dataset. In that case, we argue that we should also use essays’ string length and word count as quality indicators for the available data. Other contexts that may be useful could be the counts of EDUs and ACs identified in an essay to further act as an indicator of quality for the model.
@@ 2.1 Background
As discussed before, AES has emerged as a prominent field, with specific critical contributions since the 1950s shaping its development. Ellis Batten Page’s work has been widely acknowledged as a fundamental pillar in the advancement of AES [31]. As early as 1966, Page advocated for the feasibility of computer-based essay scoring [29]. In 1968, he published his influential work on Project Essay Grade (PEG), a specialised programme tailored for essay evaluation [30]. Subsequently, Page collaborated with esteemed institutions like the Educational Testing Service (ETS), engaging in continuous refinement and updates to PEG, which culminated in successful trials in 1994 [32].
Another noteworthy contribution to AES came from Peter Foltz and Thomas Landauer, who introduced the Intelligent Essay Assessor (IEA) in 1997 to score essays in undergraduate courses [12]. IEA has since evolved into a product offered by Pearson Educational Technologies and has gained extensive usage across commercial applications and state and national exams [1].
Within this landscape, ETS played a pivotal role in advancing AES by developing ”e-rater,” an automated essay scoring program first employed in commercial settings in February 1999 [4]. Jill Burstein, a prominent figure, took the lead in spearheading the team responsible for its development. ETS’s Criterion Online Writing Evaluation Service leverages the e-rater engine to provide users with comprehensive scores and targeted feedback on their writing performance.
@@ 2.2 Automated Student Assessment Prize
The Hewlett Foundation organised the Automated Student Assessment Prize (ASAP) in collaboration with Kaggle in 2012 to evaluate AES systems’ reliability. It predicted human raters’ essay scores [2]. The competition attracted 201 participants and featured a demonstration with nine AES vendors on a subset of the ASAP dataset. The ASAP-AES dataset remains the critical resource for AES research [19], including our study.
However, the claim of AES equivalence to human raters has drawn scrutiny [35]. Concerns include methodology issues like paragraph-based datasets, reliance on human graders for specific datasets, and use of the ”resolved score” instead of the true score [13]. Addressing these points is essential for robustly evaluating AES systems.
Jeon and Strube [15] noted a length-score correlation in the ASAP-AES dataset except for one prompt. Correlation coefficients were high (Pearson: 0.702, Spearman: 0.707, p-values << 0.001). Perelman [35] also raised this bias in 2014.
@@ 2.3 Current Approaches
To give a brief overlay of the current approaches, we will shortly discuss some recent papers that significantly impacted the scientific discourse in AES. In 2018, Ye Tai et al. [40] introduced a novel approach, incorporating neural coherence features into an end-to-end model for text scoring. Their system outperformed existing methods on benchmark datasets, demonstrating the need for more accurate and robust approaches. In the same year, contemporary limitations were addressed by Cozma et al. [8] by utilising string kernels to extract features from essays and word embeddings to capture semantic meaning. Their approach exhibited greater flexibility and adaptability in handling diverse text types and capturing nuanced aspects of language use.
@@
Another interesting take in 2021 was that of Ma et al., who modelled essays’ hierarchical structure [23]. They proposed a novel approach that utilises a hierarchical graph structure based on Graph Convolutional Networks (GCN) to encode the document structure of essays. By leveraging graph aggregation, their method captures structured coherence and discourse information. Experimental evaluations on the ASAP dataset demonstrate the effectiveness of their approach in enhancing AES. This implied that using GCNs could benefit the AES task if the latent hierarchical structure in essays can be exploited well.
@@
A novel approach to AES came with the implementation of the concept of Multi-Task Learning (MTL) as introduced by Kumar et al. [18] for AES systems, as opposed to Single-Task Learning (STL). Their study demonstrated the superiority of MTL systems utilising Long Short-Term Memory (LSTM) and Bi-Directional LSTMs (Bi-LSTMs), outperforming STL systems. Wang et al. in 2022 [44] employed BERT in a joint learning framework, capturing multi-scale representations of essays to enhance performance. They utilised various loss functions and transfer learning techniques, including out-of-domain data, to further improve the system’s capabilities. Their work also introduced a custom loss function called Margin Ranking, discussed later in Section 4.4. This helps balance out the model during training as it provides contextual ordering of the essays present in the batch based on their score. Table 6.5 shows their best-performing model’s scores on the dataset.
@@
The limited, somewhat noisy data was one of the primary reasons, as pointed out by Wang et al. in their work which results in the sub-par performance of transformer-based models. Jong et al. [16] addressed this issue of limited diversity in training data by proposing augmentation techniques such as back translations and adjusted scores for AES. Yang et al. [46] proposed a method for fine-tuning pre-trained language models using multiple losses. Their approach combined mean square error and batch-wise ListNet losses to enhance score calculation. They called it a combination of regression and re-ranking, and their model was thus called $R^{2}$-BERT as noted in Table 6.5. Theirs was also the model that held the state-of-the-art performance before Tashu et al.’s approach was implemented.
@@
More recently, Tashu et al. [39] introduced an innovative approach by using Multi-Channel Convolutional Neural Networks (CNNs) together with Recurrent Neural Networks (RNNs). Their architecture was named DeLAES and was trained on prompt-specific data, i.e., trained separately on each essay set. This helped them model the n-gram attributes (2-grams, 3-grams and 4-grams) from the word embeddings. These were then consolidated into comprehensive essay-level vectors utilising the process of max-pooling. Additionally, the architecture incorporates the utilisation of Bidirectional Gated Recurrent Units (BGRUs), enabling access to antecedent and subsequent contextual representations. Their work established the architecture’s significance in AES as the new state-of-the-art. The performance of their model is also noted in Table 6.5.
@@ 2.4 Latent Structure
This section will explore potential ways to leverage the latent structure inherent in text-based inputs. An Elementary Discourse Unit (EDU), which is a sub-sentence phrase unit, was first proposed from Rhetorical Structure Theory (RST) authored by Mann and Thompson [24]. It offers a holistic context level in that it acts as a complete unit of meaning and thus could help a model identify a complete thought or idea. A more succinct context for our use case would be that of an Argument Component (AC). The analysis of argumentation within discourse holds a critical role in argumentation theory, and central to this analysis is the concept of ACs, which provides insight into the structural composition of arguments. Van Eemeren and Grootendorst’s influential work [42] set a significant milestone in understanding argumentation analysis. While this seminal piece of literature does not explicitly introduce the term ”Argument Component,” it provides a structured framework for identifying and assessing the discrete foundational elements that constitute arguments.
Another important note is that, in developing a model to detect ACs, it’s crucial to consider a concise study by Lawrence et al. [20]. Their research emphasises the correlation between the Rhetorical Structure Theory (RST) discourse structure and argumentative constructs. They show that the discourse relations in RST often align with argumentative relations. This sheds light on the potential to use EDU-level context to enhance a model’s ability to understand the concept of ACs better.
@@
The data required for this project involved three major datasets. ASAP-AES dataset [2] was used for the Automated Essay Scoring (AES) task, the RST Discourse Treebank dataset [5] was used for Elementary Discourse Unit (EDU) span labelling and the Argument Annotated Essays (V2) [38] was used for Argument Component (AC) span labelling.
@@ 3.1 ASAP-AES
The ASAP-AES dataset[2], developed by the Hewlett Foundation, is a pivotal resource in AES research. With diverse essays paired with human-scored responses, it facilitates the evaluation of AES algorithms. It comprises essays belonging to separate sets. This sectioning was done because each set of essays was written in response to a particular prompt or question specific to that essay set.
@@ 3.2 RST Discourse Treebank
The RST Discourse Treebank (RST-DT) [5] is a pivotal resource for discourse analysis, offering annotated text segments with discourse structure. Developed by Lynn Carlson, Daniel Marcu, and Mary Ellen Okurowski based on the RST framework of Mann and Thompson, it aids in studying text coherence and discourse relations. There is a total of 385 Wall Street Journal articles taken from the Penn Treebank divided into 347 documents in the training set and 38 in the test set. The input data consists of sentences split by a newline character, and the true output has the EDU text split by newline characters. A sample of this is as follows:
@@
Based on the original work of Wang et al. [43], we randomly sampled 34 articles (roughly 10%) from the train set. Precision, Recall, and primarily F1-scores were used, as discussed in Section 5.1. Our primary interests in this dataset were the raw text and the segmented EDUs for each essay individually.
@@ 3.3 Argument Annotated Essays (version 2)
The Argument Annotated Essays (Version 2) dataset [38] is a significant resource for studying argumentation structures in essays. It offers annotated essays with labelled argument components, fostering research in argument analysis.
@@
To discuss the methodology implemented in this paper, we need to reiterate the primary research question —What can be done to provide transformers with relevant context such that it performs well in the downstream task of Automated Essay Scoring? As discussed in Chapter 1, this raises several questions. Based on the work of Mercier and Sperber’s [25], we assume that an essay is essentially a textual piece arguing in favour of some standpoint and is trying to convince the reader of the same. We also reiterate here that we restrict our scope to the following four types of context as discussed in Section 1:
@@
Our initial candidate for the transformer architecture was RoBERTa [22]. While BERT was the standard architecture for all of the discussed transformer-based models used for AES in Section 2, RoBERTa performs better than BERT across several benchmarks, as the source paper claimed. It was trained using a more extensive and diverse dataset with longer training times. The enhanced pre-training of RoBERTa helps it capture nuanced contextual relationships in language, making it well-suited for tasks that require a deep understanding of meaning and context —precisely what is needed for our use-case. It also incorporates techniques like dynamic masking and larger batch sizes during pre-training, leading to better regularisation and less overfitting. The decision to use a standard classifier head was made such that we could easily compare the performance of the baseline model with those of the context-augmented data, which involved changes to the architecture parameters that were minimal (only in the case of feature-based contexts) to none (in every other scenario). The architecture is demonstrated graphically using Figure 4.1.

### v2

@@ Acknowledgements.
However, it is important to note that the effectiveness of an AES system hinges on the underlying algorithms and models. There has been significant research using different model architectures in the AES task, and although transformer-based models have shown remarkable success in various natural language processing tasks [min2021recent], their performance in AES tasks has revealed certain limitations. This was demonstrated by Tashu et al. [tashu2022deep], in contrast to the limitations discussed by Wang et al. [wang_2022_on]. By using a Convolution Neural Network (CNN) and Recurrent Neural Network (RNN) based architecture, Tashu’s DeLAES model significantly outperformed the transformer-based state-of-the-art model of Yang et al. [yang-etal-2020-enhancing]. It is important to note that none of these transformer-based models implemented proper contextual support to leverage the data fully. This inadequacy sparks the need to explore context augmentation techniques that could potentially enhance the capabilities of transformer-based AES models.
Contextual cues have been found to impact human understanding of language significantly [FERSTL202137], and there is a growing recognition of their potential in improving AI-driven language tasks[Hung2014]. Our research aims to harness this potential by investigating the integration of context in the context of AES. At the core of this research lies a fundamental question:
@@
We assert that an essay can be described as a textual composition crafted to persuade the reader of the author’s perspective. Consequently, essays primarily strive to argue for a specific standpoint concerning the topic at hand. We substantiate this based on Mercier and Sperber’s research [mercier_sperber_2011], which underscores the intrinsic argumentative nature of human reasoning, further reinforcing the inherently persuasive quality of essays.
To restrict our scope of research, we limited our search to the following granularities: Relative Context, Prompt Context, Structural Context and Feature-based Context. The essays are graded on a curve, implying that an individual essay would have a relative rank amongst all the other essays for a specific prompt. Margin Ranking, as discussed in Section 4.4, explores this in the form of relative context. As discussed in Section 3.1, the dataset has essays collected from eight different sets where every essay is written in response to their respective prompt. Understanding the prompt or the topic should be crucial to assess the essay adequately. This should give the model an insight into the essay’s key elements, ideally better identifying unique patterns for good and bad aspects of a given prompt. Thus, a prompt can successfully act as a proper context. Essays may follow different genres and writing styles, such as argumentative, narrative, or expository. However, we assume that an essay’s primary objective is to convince the reader of some truth. To this end, if we provide the latent structural context of the essay to the model, it may understand the interdependencies of the sentences better. To address this, we want to introduce the structural contexts at the scale of Elementary Discourse Units (EDUs) [rst-edu] and Argument Components (ACs) [lawrence-reed-2019-argument]. As discussed in Section 2.2, the length is correlated with the scores in the dataset. In that case, we argue that we should also use essays’ string length and word count as quality indicators for the available data. Other contexts that may be useful could be the counts of EDUs and ACs identified in an essay to further act as an indicator of quality for the model.
@@ 2.1 Background
As discussed before, AES has emerged as a prominent field, with specific critical contributions since the 1950s shaping its development. Ellis Batten Page’s work has been widely acknowledged as a fundamental pillar in the advancement of AES [page2003project]. As early as 1966, Page advocated for the feasibility of computer-based essay scoring [page66]. In 1968, he published his influential work on Project Essay Grade (PEG), a specialised programme tailored for essay evaluation [page68]. Subsequently, Page collaborated with esteemed institutions like the Educational Testing Service (ETS), engaging in continuous refinement and updates to PEG, which culminated in successful trials in 1994 [page1994].
Another noteworthy contribution to AES came from Peter Foltz and Thomas Landauer, who introduced the Intelligent Essay Assessor (IEA) in 1997 to score essays in undergraduate courses [Foltz1999]. IEA has since evolved into a product offered by Pearson Educational Technologies and has gained extensive usage across commercial applications and state and national exams [PearsonAssessmentsK12].
Within this landscape, ETS played a pivotal role in advancing AES by developing ”e-rater,” an automated essay scoring program first employed in commercial settings in February 1999 [Burstein2003TheES]. Jill Burstein, a prominent figure, took the lead in spearheading the team responsible for its development. ETS’s Criterion Online Writing Evaluation Service leverages the e-rater engine to provide users with comprehensive scores and targeted feedback on their writing performance.
@@ 2.2 Automated Student Assessment Prize
The Hewlett Foundation organised the Automated Student Assessment Prize (ASAP) in collaboration with Kaggle in 2012 to evaluate AES systems’ reliability. It predicted human raters’ essay scores [asap-aes]. The competition attracted 201 participants and featured a demonstration with nine AES vendors on a subset of the ASAP dataset. The ASAP-AES dataset remains the critical resource for AES research [asapreview], including our study.
However, the claim of AES equivalence to human raters has drawn scrutiny [PERELMAN2014104]. Concerns include methodology issues like paragraph-based datasets, reliance on human graders for specific datasets, and use of the ”resolved score” instead of the true score [Hamner2012ContrastingSA]. Addressing these points is essential for robustly evaluating AES systems.
Jeon and Strube [jeon-strube-2021-countering] noted a length-score correlation in the ASAP-AES dataset except for one prompt. Correlation coefficients were high (Pearson: 0.702, Spearman: 0.707, p-values << 0.001). Perelman [PERELMAN2014104] also raised this bias in 2014.
@@ 2.3 Current Approaches
To give a brief overlay of the current approaches, we will shortly discuss some recent papers that significantly impacted the scientific discourse in AES. In 2018, Ye Tai et al. [tay_2018_skipflow] introduced a novel approach, incorporating neural coherence features into an end-to-end model for text scoring. Their system outperformed existing methods on benchmark datasets, demonstrating the need for more accurate and robust approaches. In the same year, contemporary limitations were addressed by Cozma et al. [cozma-etal-2018-automated] by utilising string kernels to extract features from essays and word embeddings to capture semantic meaning. Their approach exhibited greater flexibility and adaptability in handling diverse text types and capturing nuanced aspects of language use.
@@
Another interesting take in 2021 was that of Ma et al., who modelled essays’ hierarchical structure [Ma2021]. They proposed a novel approach that utilises a hierarchical graph structure based on Graph Convolutional Networks (GCN) to encode the document structure of essays. By leveraging graph aggregation, their method captures structured coherence and discourse information. Experimental evaluations on the ASAP dataset demonstrate the effectiveness of their approach in enhancing AES. This implied that using GCNs could benefit the AES task if the latent hierarchical structure in essays can be exploited well.
@@
A novel approach to AES came with the implementation of the concept of Multi-Task Learning (MTL) as introduced by Kumar et al. [kumar-etal-2022-many] for AES systems, as opposed to Single-Task Learning (STL). Their study demonstrated the superiority of MTL systems utilising Long Short-Term Memory (LSTM) and Bi-Directional LSTMs (Bi-LSTMs), outperforming STL systems. Wang et al. in 2022 [wang_2022_on] employed BERT in a joint learning framework, capturing multi-scale representations of essays to enhance performance. They utilised various loss functions and transfer learning techniques, including out-of-domain data, to further improve the system’s capabilities. Their work also introduced a custom loss function called Margin Ranking, discussed later in Section 4.4. This helps balance out the model during training as it provides contextual ordering of the essays present in the batch based on their score. Table 6.5 shows their best-performing model’s scores on the dataset.
@@
The limited, somewhat noisy data was one of the primary reasons, as pointed out by Wang et al. in their work which results in the sub-par performance of transformer-based models. Jong et al. [jong2022improving] addressed this issue of limited diversity in training data by proposing augmentation techniques such as back translations and adjusted scores for AES. Yang et al. [yang-etal-2020-enhancing] proposed a method for fine-tuning pre-trained language models using multiple losses. Their approach combined mean square error and batch-wise ListNet losses to enhance score calculation. They called it a combination of regression and re-ranking, and their model was thus called $R^{2}$-BERT as noted in Table 6.5. Theirs was also the model that held the state-of-the-art performance before Tashu et al.’s approach was implemented.
@@
More recently, Tashu et al. [tashu2022deep] introduced an innovative approach by using Multi-Channel Convolutional Neural Networks (CNNs) together with Recurrent Neural Networks (RNNs). Their architecture was named DeLAES and was trained on prompt-specific data, i.e., trained separately on each essay set. This helped them model the n-gram attributes (2-grams, 3-grams and 4-grams) from the word embeddings. These were then consolidated into comprehensive essay-level vectors utilising the process of max-pooling. Additionally, the architecture incorporates the utilisation of Bidirectional Gated Recurrent Units (BGRUs), enabling access to antecedent and subsequent contextual representations. Their work established the architecture’s significance in AES as the new state-of-the-art. The performance of their model is also noted in Table 6.5.
@@ 2.4 Latent Structure
This section will explore potential ways to leverage the latent structure inherent in text-based inputs. An Elementary Discourse Unit (EDU), which is a sub-sentence phrase unit, was first proposed from Rhetorical Structure Theory (RST) authored by Mann and Thompson [rst-edu]. It offers a holistic context level in that it acts as a complete unit of meaning and thus could help a model identify a complete thought or idea. A more succinct context for our use case would be that of an Argument Component (AC). The analysis of argumentation within discourse holds a critical role in argumentation theory, and central to this analysis is the concept of ACs, which provides insight into the structural composition of arguments. Van Eemeren and Grootendorst’s influential work [pdafr] set a significant milestone in understanding argumentation analysis. While this seminal piece of literature does not explicitly introduce the term ”Argument Component,” it provides a structured framework for identifying and assessing the discrete foundational elements that constitute arguments.
Another important note is that, in developing a model to detect ACs, it’s crucial to consider a concise study by Lawrence et al. [lawrence-reed-2019-argument]. Their research emphasises the correlation between the Rhetorical Structure Theory (RST) discourse structure and argumentative constructs. They show that the discourse relations in RST often align with argumentative relations. This sheds light on the potential to use EDU-level context to enhance a model’s ability to understand the concept of ACs better.
@@
The data required for this project involved three major datasets. ASAP-AES dataset [asap-aes] was used for the Automated Essay Scoring (AES) task, the RST Discourse Treebank dataset [carlson2002rst] was used for Elementary Discourse Unit (EDU) span labelling and the Argument Annotated Essays (V2) [aae2] was used for Argument Component (AC) span labelling.
@@ 3.1 ASAP-AES
The ASAP-AES dataset[asap-aes], developed by the Hewlett Foundation, is a pivotal resource in AES research. With diverse essays paired with human-scored responses, it facilitates the evaluation of AES algorithms. It comprises essays belonging to separate sets. This sectioning was done because each set of essays was written in response to a particular prompt or question specific to that essay set.
@@ 3.2 RST Discourse Treebank
The RST Discourse Treebank (RST-DT) [carlson2002rst] is a pivotal resource for discourse analysis, offering annotated text segments with discourse structure. Developed by Lynn Carlson, Daniel Marcu, and Mary Ellen Okurowski based on the RST framework of Mann and Thompson, it aids in studying text coherence and discourse relations. There is a total of 385 Wall Street Journal articles taken from the Penn Treebank divided into 347 documents in the training set and 38 in the test set. The input data consists of sentences split by a newline character, and the true output has the EDU text split by newline characters. A sample of this is as follows:
@@
Based on the original work of Wang et al. [wang-etal-2018-toward], we randomly sampled 34 articles (roughly 10%) from the train set. Precision, Recall, and primarily F1-scores were used, as discussed in Section 5.1. Our primary interests in this dataset were the raw text and the segmented EDUs for each essay individually.
@@ 3.3 Argument Annotated Essays (version 2)
The Argument Annotated Essays (Version 2) dataset [aae2] is a significant resource for studying argumentation structures in essays. It offers annotated essays with labelled argument components, fostering research in argument analysis.
@@
To discuss the methodology implemented in this paper, we need to reiterate the primary research question —What can be done to provide transformers with relevant context such that it performs well in the downstream task of Automated Essay Scoring? As discussed in Chapter 1, this raises several questions. Based on the work of Mercier and Sperber’s [mercier_sperber_2011], we assume that an essay is essentially a textual piece arguing in favour of some standpoint and is trying to convince the reader of the same. We also reiterate here that we restrict our scope to the following four types of context as discussed in Section 1:
@@
Our initial candidate for the transformer architecture was RoBERTa [liu2019roberta]. While BERT was the standard architecture for all of the discussed transformer-based models used for AES in Section 2, RoBERTa performs better than BERT across several benchmarks, as the source paper claimed. It was trained using a more extensive and diverse dataset with longer training times. The enhanced pre-training of RoBERTa helps it capture nuanced contextual relationships in language, making it well-suited for tasks that require a deep understanding of meaning and context —precisely what is needed for our use-case. It also incorporates techniques like dynamic masking and larger batch sizes during pre-training, leading to better regularisation and less overfitting. The decision to use a standard classifier head was made such that we could easily compare the performance of the baseline model with those of the context-augmented data, which involved changes to the architecture parameters that were minimal (only in the case of feature-based contexts) to none (in every other scenario). The architecture is demonstrated graphically using Figure 4.1.